Suppose that the labour supply curve for a large university in a small town is given by w = 40+0.07L where L is number of units of labour per week and w is the weekly wage paid per unit of labour. If the university is currently hiring 1000 units of labour per week, the marginal cost of an additional unit of labour.


A. equals the wage rate

B. equals the wage rate plus 210

C. equals the wage rate plus 70.

D. equals the wage rate plus 140


1
Expert's answer
2020-12-23T07:31:26-0500

D. equals the wage rate plus 140


w = 40+0.07L


W = 40+ ((0.07)(1000))


W = 40 +70


W = 110 +30


W = 140

Therefore:


D. equals the wage rate plus 140